OpenStudy (anonymous):

solve the following system of equations 3x+3y+2z=-5 , X + Y + 3X = -4 , 2Y +2Y+5z =-7

OpenStudy (blurbendy):

youll need to use the elimination method

OpenStudy (anonymous):

are u sure that the question is correct

OpenStudy (anonymous):

yeah i just forgot to put 3x+3y+2z=-5 On The Next Line

hero (hero):

@sammiee10, in equations two and three, some of your variables repeat

hero (hero):

Are you going to fix that? Or did you want to continue with using 2y + 2y + 5z = -7 4y + 5z = -7